OpenStudy (anonymous):

HELP MEDALS+FAN!!!! Evaluate. 8 • [5 + (18 – 7)] ÷ 4

OpenStudy (anonymous):

plz watch as i will simplify this quickly 8*(5+11)/4 8*16/4 8*4

OpenStudy (anonymous):

anymore questions?

OpenStudy (anonymous):

yes Which property justifies this statement? 2 + (8 + 9) = (2 + 8) + 9 A. transitive B. commutative C. distributive D. associative

OpenStudy (anonymous):

transitive: a=b, b=c, so a=c commutative: a+b=b+a distributive: a(b+c)=ab+ac associative: a+(b+c)=(a+b)+c

OpenStudy (anonymous):

thank u now Translate the word phrase into a math expression. nine more than five times a number A. 9 > 5n B. 9 – 5n C. 5n – 9 D. 5n + 9

OpenStudy (anonymous):

9 more than 5 times a # 9+5 times a number 9+5*x

OpenStudy (anonymous):

well, in your case its n not x

OpenStudy (anonymous):

thanxs 1 more Which situation could be solved using this expression? 4k + 2 A. Jason read four more than twice the number of books that Marty read. How many books did Marty read? B. Jason read half as many books as Marty read. How many books did Jason and Marty both read? C. Jason read two more than four times the number of books that Marty read. How many books did Jason read? D. Jason read four times as many books as Marty. Marty read two books. How many books did they read together?

OpenStudy (anonymous):

4 time a number +2 2 more than 4 times a number look for those key words above this statement

OpenStudy (anonymous):

the key words were 2 more than 4 times a number if that was confusing

OpenStudy (anonymous):

thank u so much